Calculate the pH of each of the following strong acid solutions. (a) 0.00594 M HBrO4 pH = (b) 0.885 g of HBrO4 in 41.0 L of solution pH = (c) 20.0 mL of 8.90 M HBrO4 diluted to 3.80 L pH = (d) a mixture formed by adding 44.0 mL of 0.00256 M HBrO4 to 79.0 mL of 0.00425 M HBrO4 pH =
Answer #1

(a)

[H+] = 0.00594 M

pH = -log (H+)

     = -log (0.00594)

pH = 2.23

(b)

mass of HBrO4 = 0.885 g

moles of HBrO4 = 0.885 / 144.91 = 6.11 x 10^-3 mol

concentration = 6.11 x 10^-3 / 41.0 = 1.50 x 10^-4 M

pH = -log (1.50 x 10^-4)

pH = 3.83

(c)

20.0 x 8.90 = 3800 x M

M = 0.0468 M

pH = -log (0.0468)

pH = 1.33

(d)

[H+] = 44 x 0.00256 + 79 x 0.00425 / (44 + 79)

       = 3.64 x 10^-3 M

pH = 2.44